if the perimeter of a rectangle is 112 meter, and the length is one meter shorter than twice the width, then what is the length and the width

Answer by asuar010(338)

the perimeter formula for a rectangle is P=2(l+w) l=lenght, w=width
by interpreting the problem we assign x as the width and 2x-1 as the length...
w=x
l=2x-1
by substituting the values we get:
112=2(x+2x-1) then 112=2(3x-1) then 112=6x-2 then 114=6x and we obtain x=19.
then by subtituting the value of x we get that w=19 and l=2(19)-1=37